OpenStudy (anonymous):

How many photons are in a laser pulse of 0.363J at 657 nm?

OpenStudy (anonymous):

E=h*v \[E=(h*c)/\] E=((6.62*10^-34)*(3*10^8))/(657*10^-9) E=3.025*10^-19 J, this is the energy of one photon N=Np/Nph N=(.363)/(3.025*10^-19) N=119.9*10^16 photons